# Subgraph-Oriented Testing for Deep Learning Library

### v1

@@ V-A RQ1: Validity of Input
To understand the effectiveness of SORT in generating valid inputs, we compared the ratio of valid inputs generated by different methods for all 15 target APIs listed in Section IV-E. For a fair comparison, we run each method to generate a total of 15,000 valid test inputs across 15 target APIs. For baseline methods, we tried to generate 1,000 valid test inputs for each API by following their setup and we recorded the total number of test inputs generated [^4] In cases where generating 1,000 valid test cases required more than 10,000 attempts, we halted further testing for cost efficiency.. For SORT, we generate a total of 15,000 valid test inputs for APIs in all subgraphs.
@@
DocTer shows an 11.00% ratio of valid inputs. Its ineffectiveness results from imprecise input constraints extracted from incomplete documents [16, 21]. ACETest-C and ACETest-G [^7] ACETest-C executes testing on CPU and ACETets-G executes the testing on GPU. show ratios of valid inputs of 6.80% and 6.72%. They generate invalid inputs due to their inability to accurately solve constraints in input validation loops, C++ pointers, and tensor constraints [21]. When they fail to extract precise constraints, they will randomly generate values, including excessive values, as inputs for APIs. In comparison, SORT avoids this problem by generating input values based on the features of runtime input data, which are more likely to be valid and better simulate real-life inputs. For example, as shown in Fig. 11, ACETest generates a test input with an uncommonly large dropout probability, resulting in a runtime error: “dropout probability has to be between 0 and 1”. Meanwhile, SORT generates the parameter p as 0.5 for dropout based on the runtime data value and does not trigger runtime errors.
@@
To summarize, SORT achieves a higher ratio of valid inputs.
V-B RQ2: Effectiveness of Finding Precision Bugs
@@
Meanwhile, all baseline methods missed these precision bugs. They generally focused on crash and performance bugs. For instance, FreeFuzz detected 267 performance bugs, primarily in the adaptive_avg_pool2d API (241 cases). These bugs were associated with unexpected performance behavior, where the speed of processing float32 values outperformed float16 values. DeepREL identified 7 crash bugs across __mul__, div, and __add__, typically related to input types and boundary cases. DocTer uncovered 236 crash bugs, primarily in max_pool2d (219 cases) and conv2d (17 cases), caused by corner cases. TitanFuzz found 223 crash bugs, notably in matmul (75 cases) and adaptive_avg_pool2d (46 cases), as well as 30 NaN bugs, with a concentration in gelu (21 cases) and matmul (9 cases). Finally, ACETest did not detect any bugs during our experiments. As a reminder, when reproducing FreeFuzz, DeepREL, DocTer, TitanFuzz, and ACETest methods, we observed a decrease in their error-triggering ability compared to the results reported in their papers. This can be due to the newer version of PyTorch (2.3.0) used in our experiment, where many bugs have been fixed.
In summary, SORT outperformed baseline methods in detecting precision bugs.
V-C RQ3: Contribution of Interaction for Bug-Detection Ability
@@
In this section, we investigate whether interaction can contribute to the bug-detection ability of SORT. First, we compare the average output differences observed between CPU and GPU executions for each API triggered by SORT and single-API testing. Fig. 13(a) shows the comparison, where the red dotted line indicates the threshold to report a bug. Differences below the threshold are considered tolerant precision differences. We found that flatten, __mul__, batch_norm, dropout, relu, gelu, and __add__ show average observed differences greater than thresholds in SORT but not in single-API testing. This trend suggests that SORT can magnify the differences between CPU and GPU executions and thereby is more likely to reveal more bugs in APIs.
@@
Next, we further analyze how API interactions help to reveal bugs by magnifying output differences. Given that we cannot confirm the buggy APIs, we do not study the direct relation between API interactions and actual bug detection ability in this work. Meanwhile, it is well-known that only if an API is involved in an observed abnormal behavior, can the bugs of this API be revealed based on the observed behavior. Thus, we study how API interactions help to trigger abnormal behaviors involving more APIs, thereby likely contributing to revealing more buggy APIs. Specifically, we study the frequency of each API being involved in erroneous behaviors observed in test executions. This frequency is defined as the ratio of the test executions where an API is involved in erroneous behavior to all executions in which this API is executed. If an API is more frequently involved in observed abnormal behaviors, developers may pay more attention to inspecting this API with the observed behaviors as clues. As mentioned in Section III-D, when SORT reveals a bug, both the API whose outputs violate the difference threshold and its previous APIs in the frequent subgraph relate to the observed abnormal behavior; meanwhile, single-API testing only associates the API under testing with the observed abnormal behavior.
@@ VI Discussion
To understand whether SORT aligns with the expectations of developers on testing methods of deep learning libraries, we conducted a survey with ten developers working with DL libraries. Seven of these participants are developers from ICT companies working with DL libraries, and the other three are PhD students with over five years of DL model-building and debugging experience. The main findings of the survey are as follows:
Model-level testing leads to difficulties in error localization and testing API sequences may alleviate this problem. All participants agreed that it is difficult for developers to trace the root cause (i.e., specific buggy APIs) of errors based on the model-level testing results. They also suggest this hinders the efficiency of debugging and repair practices. Meanwhile, 60% participants believed that bugs identified through testing API sequences of no more than ten APIs are easier to trace and fix. As introduced in Section II-B, the frequent subgraphs mined by SORT typically include fewer nodes (e.g., 2 to 7 on our subjects of popular DL models), making them easier to debug. Therefore, participants’ feedback suggests that subgraph-level testing is likely to offer benefits in alleviating error localization problems, as mentioned in Section II-B.
@@
Minor precision differences below the warning threshold may accumulate and finally cause serious issues. 80% participants noted that DL library APIs often exhibit precision differences when executed on different hardware platforms and these differences can accumulate over time. They agree that accumulated differences may eventually lead to issues like non-convergence on certain hardware. The feedback from participants emphasizes the importance of identifying these minor differences, which reveals the significance of subgraph testing since it can reveal minor differences magnified by API interactions as shown in Section V-C.
VII Related Works
@@
Considering the potential helpfulness of API interactions for bug detection, later methods create API combinations [24, 25, 20]. They focus on achieving higher API combination diversity. They create a whole model as the test subject as the model-level testing methods do and thereby also suffer from difficulties in fault localization. Besides, EAGLE [26] employs equivalence rules for testing. Some of its equivalence rules involve API combinations, but no errors were detected with such combinations. In addition, TitanFuzz [18] uses large language models to generate API combinations, while many reported bugs are only triggered by the specific API sequences artificially made, which are rarely used in real-world DL models. These methods do not take care of the real API interactions that tend to be more important since they are more likely to be frequently used in daily life and thereby their bugs tend to lead to more severe consequences.
In this paper, we introduce a novel testing approach that takes the test subjects at a granularity distinct from existing methods, i.e., frequent subgraphs of DL model computation graphs. This granularity introduces meaningful API interactions while maintaining the efficiency in fault localization, and finally helps to detect more precision bugs. Besides, SORT enhances the generation of test inputs. Unlike existing works that rely on constraints extracted from static documents and codes [16, 21], SORT generates test inputs by referring to the real runtime inputs features on easily obtained datasets to simulate real valid inputs. Also, unlike existing methods overlooking the value ranges of generated inputs [14, 17, 18], SORT collects the value ranges of real inputs and generates inputs within these ranges to better simulate real executions, expecting to reveal faults likely to happen in real-life usage of DL libraries.
VIII Conclusion
In this paper, we propose a novel DL library testing method called SORT to address the limitations of existing methods in leveraging representative API interactions and preparing meaningful test inputs to test DL libraries. Specifically, to reflect meaningful API interactions while maintaining fault localization efficiency, SORT uses a new granularity of test subjects, i.e., frequent subgraphs of model computation graphs. Besides, SORT collects extensive runtime input features of each API under executing real-life data and generates test inputs likely more valid and meaningful based on these features. Evaluation results demonstrate SORT effectively reveals many unaware precision issues on popular PyTorch APIs, by generating much more (100%) valid inputs and properly magnifying minor yet impactful output differences via meaningful API interactions.

### v2

@@ V-A RQ1: Validity of Input
To understand the effectiveness of SORT in generating valid inputs, we compared the ratio of valid inputs generated by different methods for all 15 target APIs listed in Section IV-E. For a fair comparison, we run each method to generate a total of 15,000 valid test inputs across 15 target APIs. For baseline methods, we tried to generate 1,000 valid test inputs for each API by following their setup and we recorded the total number of test inputs generated [^4] In cases where generating 1,000 valid test cases required more than 10,000 attempts, we halted further testing for cost efficiency.. For SORT, we ensure a total of 15,000 valid test inputs for APIs in all subgraphs.
@@
DocTer shows an 11.00% ratio of valid inputs. Its ineffectiveness results from imprecise input constraints extracted from incomplete documents [16, 19]. ACETest-C and ACETest-G [^7] ACETest-C executes testing on CPU and ACETets-G executes the testing on GPU. show ratios of valid inputs of 6.80% and 6.72%, respectively. They generate invalid inputs due to their inability to accurately solve constraints in input validation loops, C++ pointers, and tensor constraints [19]. When they fail to extract precise constraints, they will randomly generate values, including excessive values, as inputs for APIs. In comparison, SORT avoids this problem by generating input values based on the features of runtime input data, which are more likely to be valid and better simulate real-life inputs. For example, as shown in Fig. 11, ACETest generates a test input with an uncommonly large dropout probability, resulting in a runtime error: “dropout probability has to be between 0 and 1”. Meanwhile, SORT generates the parameter p as 0.5 for dropout based on the runtime data value and does not trigger runtime errors.
@@
To summarize, by referring to the extensive features of type, shape, and value range of the runtime inputs for DL library APIs on real datasets, SORT successfully generates more valid test inputs than the baseline methods. Such generated inputs are expected to simulate real execution data of DL libraries and help to concentrate on detecting the errors more likely to happen in real-life usage of DL libraries. In RQ2, we will discuss the bugs revealed with such inputs.
V-B RQ2: Effectiveness of Detecting Precision Bugs
@@
Meanwhile, all baseline methods missed these precision bugs. They generally focused on crash and performance degradation bugs. For instance, FreeFuzz detected 267 performance degradation bugs, primarily in the adaptive_avg_pool2d API (241 cases). These bugs were associated with unexpected performance behavior, where the speed of processing float32 values outperformed float16 values. DeepREL identified 7 crash bugs across __mul__, div, and __add__, typically related to input types and boundary cases. DocTer uncovered 236 crash bugs, primarily in max_pool2d (219 cases) and conv2d (17 cases), caused by corner cases. TitanFuzz found 223 crash bugs, notably in matmul (75 cases) and adaptive_avg_pool2d (46 cases), as well as 30 NaN bugs, with a concentration in gelu (21 cases) and matmul (9 cases). Finally, ACETest did not detect any bugs during our experiments. As a reminder, when reproducing FreeFuzz, DeepREL, DocTer, TitanFuzz, and ACETest methods, we observed a decrease in their error-triggering ability compared to the results reported in their papers. This can be due to the newer version of PyTorch (2.3.0) used in our experiment, where many bugs have been fixed.
In summary, the testing results demonstrate that SORT can detect precision bugs, which are also considered a significant issue threatening the reliability of building and deployment of DL models, while the baseline methods barely detect such bugs. As discussed in RQ1, the improved validity and potentially higher similarity to the real-life inputs of the generated inputs contribute to this superiority. Besides, the consideration of API interactions further helps to expose precision errors as well, which we will discuss in RQ3.
V-C RQ3: Effectiveness of API Interactions in Bug Detection
@@
In this subsection, we investigate whether considering API interactions contributes to the bug-detection ability of SORT. First, we compare the average output differences observed between CPU and GPU executions for each API triggered by SORT and single-API testing. Fig. 13(a) shows the comparison, where the red dotted line indicates the threshold to report a bug. Differences below the threshold are considered tolerant precision differences. We found that flatten, __mul__, batch_norm, dropout, relu, gelu, and __add__ show average observed differences greater than thresholds in SORT but not in single-API testing. This trend suggests that SORT can magnify the differences between CPU and GPU executions and thereby is more likely to reveal more bugs in APIs.
@@
Next, we further analyze how API interactions help to reveal bugs by magnifying output differences. Given that we cannot confirm the buggy APIs, we do not study the explicit relation between API interactions and actual bug detection ability in this work. Meanwhile, it is well-known that only if an API is involved in an observed abnormal behavior, can the bugs of this API be revealed based on the observed behavior. Thus, we study how API interactions help to trigger abnormal behaviors involving more APIs, thereby likely contributing to revealing more buggy APIs. Specifically, we study the frequency of each API being involved in erroneous behaviors observed in test executions. This frequency is defined as the ratio of the test executions where an API is involved in erroneous behavior to all executions in which this API is involved. If an API is more frequently involved in observed abnormal behaviors, developers may pay more attention to inspecting this API with the observed behaviors as clues [41]. As mentioned in Section III-D, when SORT reveals a bug, both the API whose outputs violate the difference threshold and its previous APIs in the frequent subgraph relate to the observed abnormal behavior; meanwhile, single-API testing only associates the API under testing with the observed abnormal behavior.
@@ VI Discussion
To understand whether SORT aligns with developers’ expectations of testing methods of deep learning libraries, we conducted a survey with ten developers working with DL libraries. The questions and received feedback of the survey are available in our artifact [27]. Seven of these participants are developers from ICT companies working with DL libraries, and the other three are PhD students with over five years of DL model-building and debugging experience. The main findings of the survey are as follows:
Model-level testing leads to difficulties in error localization and testing API sequences may alleviate this problem. All participants agreed that it is difficult for developers to trace the root cause (i.e., specific buggy APIs) of errors based on the model-level testing results. They also suggest this hinders the efficiency of debugging and repair practices. Meanwhile, 60% participants believed that bugs identified through testing API sequences of no more than ten APIs are easier to trace and fix. As introduced in Section II-B, the frequent subgraphs mined by SORT typically include fewer nodes (e.g., 2 to 7 on our subjects of popular DL models), making them easier to debug. Therefore, participants’ feedback suggests that frequent subgraph-level testing is likely to offer benefits in alleviating error localization problems, as mentioned in Section II-B.
@@
Minor precision differences below the warning threshold may accumulate and finally cause serious issues. 80% participants noted that DL library APIs often exhibit precision differences when executed on different hardware platforms and these differences can accumulate over time. They agreed that accumulated differences may eventually lead to issues like non-convergence on certain hardware. The feedback from participants emphasizes the importance of identifying these minor differences, which reveals the significance of subgraph testing since it can reveal minor differences magnified by API interactions as shown in Section V-C.
VII Threats to Validity
A threat to validity of our approach is about the completeness of input features recorded based on certain real datasets. Indeed, we cannot guarantee to obtain the complete features of all valid inputs for each API. As a result, we may fail to generate valid inputs for the real usage scenarios not covered by the adopted datasets. However, unlike the limitations of existing methods, such incompleteness does not lead to the generation of invalid test inputs and the miss detection of non-crash bugs. Meanwhile, we have identified a few precision bugs using the information available in our evaluation. Therefore, we consider the evaluation results should demonstrate the effectiveness of our input generation approach. Besides, given the diversity of available DL datasets, supplementing the recorded features with more datasets should not be too challenging.
Another potential threat is about the selection of parameters in frequent subgraph extraction. The choice of the frequency threshold decides the frequent subgraphs available for testing DL library APIs. A too-large threshold may confine testable APIs to only the ones in highly pervasive interaction patterns; meanwhile, prioritizing the tests with patterns at a small frequency (such patterns are likely quite massive) may postpone revealing the bugs more likely to be encountered. In this work, we consider frequent subgraphs present at least five times in the collected models, representing the top 20% popular API interaction patterns. The users of SORT may follow our practice to set a proper threshold for their test needs.
The last potential threat is about the generalizability of our approach. In this work, we evaluate SORT by testing several PyTorch APIs using the popular DL models implemented by OpenMMLab. We use models in OpenMMLab because it is a popular open-source DL model zoo, and we only test PyTorch since OpenMMLab mainly provides PyTorch implementations of models. The idea of SORT is not restricted to these subjects. It can also be applied to testing other DL libraries based on the DL models implemented with those libraries from other authentic sources. One of our future works would be applying SORT to reveal bugs on other widely-used DL libraries with other DL models and datasets.
VIII Related Works
@@
Considering the potential helpfulness of API interactions for bug detection, later methods create API combinations [23, 24, 25]. Prior works focus on achieving higher API combination diversity. They create a whole model as the test subject as the model-level testing methods do and thereby also suffer from difficulties in fault localization. Besides, EAGLE [26] employs equivalence rules for testing. Some of its equivalence rules involve API combinations, but no errors were detected with such combinations. In addition, TitanFuzz [18] uses large language models to generate API combinations, while many reported bugs are only triggered by the specific API sequences artificially made, which are rarely used in real-world DL models. These methods do not take care of the real API interactions that tend to be more important since they are more likely to be frequently used in daily life and thereby their bugs tend to lead to more severe consequences.
In this paper, we introduce a novel DL library testing method, SORT, which takes the test subjects at a granularity distinct from existing methods, i.e., frequent subgraphs of DL model computation graphs. Frequent subgraph-oriented testing introduces meaningful API interactions while maintaining the efficiency in fault localization, and finally helps to detect more precision bugs. Besides, SORT enhances the generation of test inputs. Unlike existing works that rely on constraints extracted from static documents and codes [16, 19], SORT generates test inputs by referring to the real runtime inputs features on easily obtained real DL datasets to simulate real valid inputs. Also, unlike existing methods overlooking the value ranges of generated inputs [14, 17, 18], SORT collects extensive input features about the type, shape, and value ranges of real inputs and generates inputs within these ranges to better simulate real executions, expecting to reveal faults likely to happen in real-life usage of DL libraries.
IX Conclusion
In this paper, we propose a novel DL library testing method called SORT to address the limitations of existing methods in leveraging representative API interactions and preparing meaningful test inputs to test DL libraries. Specifically, to reflect meaningful API interactions while maintaining fault localization efficiency, SORT uses a new type of test subjects, i.e., frequent subgraphs of model computation graphs. Besides, SORT collects extensive runtime input features of each API under executing real-life data and generates test inputs likely more valid and meaningful based on these features. Evaluation results demonstrate SORT effectively reveals many unaware precision issues on popular PyTorch APIs, by generating much more (100%) valid inputs and properly magnifying minor yet impactful output differences via meaningful API interactions.
Acknowledgment
This work was partially supported by National Key R&D Plan of China (Grant No.2024YFF0908003), National Natural
Science Foundation of China (No. 62472326), and CCF-Zhipu Large Model Innovation Fund (No. CCF-Zhipu202408).